# Cyclic Representations of $U_{q}(\widehat{\mathfrak{sl}}_{2})$ and its Borel Subalgebras at Roots of Unity and Q-operators

### v4

@@ 2.2. Cyclic Representations of $U_{q}(\mathfrak{b}_{+})$
$\bar{\rho}_{r}(z_{0})=\bar{\rho}_{r}(z_{1})=1$
These two representations cannot be constructed by restricting full $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ representations. We need one final, $U_{q}(\mathfrak{b}_{+})$ representation $\varphi_{c}:U_{q}(\mathfrak{b}_{+})\rightarrow\operatorname{End}(W)$, $c\in\mathbb{C}^{\times}$, defined by
@@
$r\,$$\rho_{r}\sim$$r\,$$\bar{\rho}_{r}\sim$$\varphi_{c}\sim$
@@ 3.1. Factorization
For $x=z_{i}$ the equality is clear, and for $x=t_{i}$ the relation follows from the property $[\chi,Z\otimes Z]=0$ and the expression for $c_{0}$. For $x=e_{0}$, the relation becomes
@@
$r$$s$$r$$s$${\mathcal{O}}\sim$$:\Omega_{rs}\otimes\varphi_{c_{0}}\rightarrow\rho_{r}\otimes\bar{\rho}_{s}$$r$$s$$r$$s$${\mathcal{O}}^{-1}\sim$$:\rho_{r}\otimes\bar{\rho}_{s}\rightarrow\Omega_{rs}\otimes\varphi_{c_{0}}$
@@
The invertibility of ${\mathcal{P}}(Z)$ follows by a very similar argument to the proof given in Appendix A.1.
@@
$(i)\quad\mathfrak{T}_{rs}$ ::: $\rho_{r}\otimes\bar{\rho}_{s}\rightarrow\rho_{s}\otimes\bar{\rho}_{r},$
$(ii)\quad{\mathfrak{S}}_{rs}$ ::: $\bar{\rho}_{r}\otimes\rho_{s}\rightarrow\bar{\rho}_{s}\otimes\rho_{r}.$
@@ 3.2. Fusion
(3.3) $\pi_{z}(e_{1})=\begin{pmatrix}0&z\\ 0&0\end{pmatrix},\quad$ $\pi_{z}(f_{1})=\begin{pmatrix}0&0\\ z^{-1}&0\end{pmatrix},$ $\pi_{z}(t_{1})=\begin{pmatrix}q&0\\ 0&q^{-1}\end{pmatrix},\quad\pi_{z}(z_{1})=1,$
$\pi_{z}(e_{0})=\begin{pmatrix}0&0\\ z&0\end{pmatrix},\quad$ $\pi_{z}(f_{0})=\begin{pmatrix}0&z^{-1}\\ 0&0\end{pmatrix},$ $\pi_{z}(t_{0})=\begin{pmatrix}q^{-1}&0\\ 0&q\end{pmatrix},\quad\pi_{z}(z_{0})=1.$
@@
Let $r,s\in{\mathcal{C}}_{k}$, with $s=(x_{s},y_{s},\mu_{s})$. Defining $sq^{\pm 1}=(x_{s}q^{\pm 1},y_{s}q^{\pm 1},\mu_{s}q^{\pm 1})$, we see that we also have $sq^{\pm 1}\in{\mathcal{C}}_{k}$. Choose $z_{s}$ to satisfy $z_{s}^{2}=\kappa_{0}\kappa_{1}x_{s}y_{s}$ Let us define the following complex constants
@@
$(i)$ 00{0}${\rho_{sq}}$${\rho_{s}\otimes\pi_{z_{s}}}$${\rho_{sq^{-1}}}$00{0}$\scriptstyle{\iota_{s}}$$\scriptstyle{\tau_{s}}$ $\hbox{where}\quad\iota_{s}=\begin{pmatrix}c_{s}Z\\ XZ^{-1}\end{pmatrix},\;\tau_{s}=(-\frac{q}{c_{s}}XZ^{-1},Z),$
$(ii)$ 00{0}${\bar{\rho}_{sq}}$${\bar{\rho}_{s}\otimes\pi_{z_{s}}}$${\bar{\rho}_{sq^{-1}}}$00{0}$\scriptstyle{\bar{\iota}_{s}}$$\scriptstyle{\bar{\tau}_{s}}$ $\hbox{where}\quad\bar{\iota}_{s}=\begin{pmatrix}\bar{c}_{s}\\ XZ^{-2}\end{pmatrix},\;\bar{\tau_{s}}=(-\frac{1}{\bar{c}_{s}}XZ^{-2},1),$
@@
These relations are just checked by explicit computation. The injectivity and surjectivity are of course equivalent to Ker$(\iota_{s})$=0, Im($\tau_{s}$)=$W$. These relations are also simple to check. Finally the SES property then follows from the fact that $\tau_{s}\circ\iota_{s}=-qX+ZXZ^{-1}=0$.
@@ 3.3. Factorization and Fusion of L-operators
where $\mathbb{I}$ is the 2$\times$2 unit matrix. As ${L}_{\varphi_{c}}(z)$ is idependent of both $z$ and the constant $c$, we shall henceforth denote it as simply $L_{\varphi}$.
@@
$s\,$$r\,$$z$$\check{L}_{\Omega_{rs}}(z)\sim$$:\Omega_{rs}\otimes\pi_{z}\rightarrow\pi_{z}\otimes\Omega_{rs}$$z$$\check{L}_{\varphi}\sim$$:\varphi_{c}\otimes\pi_{z}\rightarrow\pi_{z}\otimes\varphi_{c}$$r\,$$z$$\check{L}_{\rho_{r}}(z)\sim$$:\rho_{r}\otimes\pi_{z}\rightarrow\pi_{z}\otimes\rho_{r}$$r\,$$z$$\check{L}_{\bar{\rho}_{r}}(z)\sim$$:\bar{\rho}_{r}\otimes\pi_{z}\rightarrow\pi_{z}\otimes\rho_{r}$
@@
$z\,$$r\,$$s$$\check{\bf L}_{\Omega_{rs}}(z)\sim$$:\pi_{z}\otimes\Omega_{rs}\rightarrow\Omega_{rs}\otimes\pi_{z}$
@@ 3.3.2. Fusion of L-operators
$(i)$$sq\,$$w$$z_{s}$$s$$sq\,$$w$$z_{s}$$s$$\quad\quad=C_{1}(s,w)$
$(ii)$$s\,$$w$$z_{s}$$sq^{-1}$$s\,$$w$$z_{s}$$sq^{-1}$$\quad\quad=C_{2}(s,w)$
$(iii)$$sq\,$$w$$z_{s}$$s$$sq\,$$w$$z_{s}$$s$$\quad\quad=C_{1}(s,w)$
$(iv)$$s\,$$w$$z_{s}$$sq^{-1}$$s\,$$w$$z_{s}$$sq^{-1}$$\quad\quad=C_{2}(s,w)$
@@
where $\check{\bf L}_{\Omega_{ss^{\prime}}}(z)$ is the $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ intertwiner $:\pi_{z}\otimes\Omega_{ss^{\prime}}\rightarrow\Omega_{ss^{\prime}}\otimes\pi_{z}$ defined above, and the coefficients are given by
@@
$(i)$$r$$r^{\prime}q$$s$$s^{\prime}$$r$$s$$s^{\prime}$$r^{\prime}$$z_{r^{\prime}}$$\quad\quad=E_{1}(r^{\prime},s,s^{\prime})$$r$$r^{\prime}q$$s$$s^{\prime}$$r$$s$$s^{\prime}$$r^{\prime}$$z_{r^{\prime}}$$(ii)$$r\,$$r^{\prime}$$s$$s^{\prime}$$r$$s$$s^{\prime}$$r^{\prime}q^{-1}$$z_{r^{\prime}}$$\quad\quad=E_{2}(r^{\prime},s,s^{\prime})$$r\,$$r^{\prime}$$s$$s^{\prime}$$r$$s$$s^{\prime}$$r^{\prime}q^{-1}$$z_{r^{\prime}}$
@@ 4. Transfer Matrices and Q-operators
4.1. Quantum Space $V^{\otimes M}$
@@
$w$$w$$w$$w$$w$$w$$Q_{\rho_{r}}(w)\sim$$r$
Let us now discuss the reason for including the twist parameter $\alpha$ in this and other definitions. First, we recall that in the case of generic $q$, we are forced to introduced such a twist in order to regularise the infinite-dimensional trace involved in the definition of the Q-operator (see for example, [VW20]). The cyclic representations in this paper are of course $N$-dimensional, and at first sight it would appear that the twist is unnecessary. However, with cyclic representations we encounter a different issue: if we choose the representation of the matrices $X,Z\in GL(N,\mathbb{C})$ to be $X_{ij}=\delta_{i,\hbox{mod}(j+1,N)}$, $Z_{ij}=\delta_{i,j}q^{i}$, for $i,j\in\{0,1,\cdots,N-1\}$, then it follows from basic properties of primitive roots of unity that we have
@@
where $\delta_{n\equiv 0}$ is defined to be equal to 1 if $n=0$ mod$(N)$, and 00 otherwise. Requiring $n\equiv 0$ is equivalent to ‘charge conservation mod $N$’ which is to be expected for transfer matrices associated with cyclic representations. Requiring that $m=0$ however means that the corresponding trace vanishes on all spin sectors apart from 00 mod $N$. The twisted trace is introduced in order to stop this phenomenon. Using the definition $\rho_{r}(t_{1})=q\mu_{r}Z^{2}$, we see that, provided we choose $2\alpha\notin\mathbb{Z}$, we have
@@
[…]
@@
$w$$w$$w$$w$$w$$w$$s$$r$$T_{\Omega_{rs}}(w):=\hbox{Tr}_{W}\left((\Omega_{rs}(t_{1}^{\alpha})\otimes\mathbb{I})L_{\Omega_{rs}}^{M}(w)\cdots L_{\Omega_{rs}}^{2}(w)L_{\Omega_{rs}}^{1}(w)\right)$
@@
$Q(z,\mu):$ $=$ $\operatorname*{Tr}_{W}({L}^{M}(z,\mu)\cdots{L}^{2}(z,\mu){L}^{1}(z,\mu)),$
$\widebar{Q}(z,\mu):$ $=$ $\operatorname*{Tr}_{W}(\widebar{L}^{M}(z,\mu)\cdots\widebar{L}^{2}(z,\mu)\widebar{L}^{1}(z,\mu)),$
@@
Hence it then follows from Proposition 4.3 and the property $[T(z),S_{z}]=0$ that $Q(z,\mu)$ and $\widebar{Q}(z,\mu)$ satisfy the standard TQ relations [^3] The dependence of the $Q$ operator on the parameter $\mu$, the algebraic necessity of this parameter, and an explanation how the dependence may be removed is explained in [VW20]
@@
(4.6) $T_{\Omega_{rs}}(w)T(z_{s}/w)=C_{1}^{M}(s,w)T_{\Omega_{r,sq}}(w)+C_{2}^{M}(s,w)T_{\Omega_{sq^{-1}}}(w).$
@@
(which can be proved either by appealing to uniqueness of the corresponding $U_{q}(\mathfrak{b}_{+})$ intertwiner or just by checking the relation), or directly from Propositions 4.2 and 4.3 $(ii)$ plus the observation that $T_{\varphi_{c}}$ commutes with $Q_{\rho_{r}}$ and $Q_{\bar{\rho}_{s}}$. Thus we have that $T_{\Omega_{rs}}(w)$ satisfies the same relations (4.6) as the Q operator of the 6-vertex model. This fact was observed in [BS90].
4.2. Quantum Space $W^{\otimes M}$
@@
$s\,$$s^{\prime}$$s\,$$s^{\prime}$$s\,$$s^{\prime}$$s\,$$s^{\prime}$$s\,$$s^{\prime}$$s\,$$s^{\prime}$$z$${\mathcal{T}}(z)=\hbox{Tr}\left((\pi_{z}(t_{1}^{\alpha})\otimes\mathbb{I})\,{\bf L}_{\Omega_{ss^{\prime}}}^{M}(z)\cdots{\bf L}_{\Omega_{ss^{\prime}}}^{2}(z){\bf L}^{1}_{\Omega_{ss^{\prime}}}(z)\right)\quad\sim$
@@
$r\,$$r^{\prime}$$r$$r^{\prime}$$s$$s^{\prime}$$s$$s^{\prime}$$s$$s^{\prime}$$s$$s^{\prime}$$s$$s^{\prime}$$s$$s^{\prime}$$s$$s^{\prime}$$s$$s^{\prime}$$s$$s^{\prime}$$s$$s^{\prime}$$s$$s^{\prime}$$s$$s^{\prime}$${\mathcal{T}}^{CP}\sim$

### v5

@@ 2.2. Cyclic Representations of $U_{q}(\mathfrak{b}_{+})$
$\bar{\rho}_{r}(z_{0})=\bar{\rho}_{r}(z_{1})=1.$
Proposition 2.2.
$\rho_{r}$ and $\bar{\rho}_{r}$ are $U_{q}(\mathfrak{b}_{+})$ representations.
Proof.
To establish this one needs to check compatibility with the $U_{q}(\mathfrak{b}_{+})$ relations $t_{i}e_{i}t_{i}^{-1}=q^{2}e_{i}$, $t_{1-i}e_{i}t_{1-i}^{-1}=q^{-2}e_{i}$ and $e_{i},[e_{i},[e_{i},[e_{1-i}]_{q^{2}}]_{1}]_{q^{-2}}=0$, for $i\in\{0,1\}$ (where $a,b]_{x}=ab-xba$). This compatibilty follows directly from the $ZX=qXZ$ and $[X,X]=0$. ∎
We need one final, $U_{q}(\mathfrak{b}_{+})$ representation $\varphi_{c}:U_{q}(\mathfrak{b}_{+})\rightarrow\operatorname{End}(W)$, $c\in\mathbb{C}^{\times}$, defined by
@@
$r\rho_{r}\simrr\,$$\bar{\rho}_{r}\sim$$\varphi_{c}\sim$
Proposition 2.3.
The $U_{q}(\mathfrak{b}_{+})$ representations $\rho_{r},\bar{\rho}_{r}$ are not restrictions of $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ representations.
Proof.
Let us proceed by contradiction. We suppose there exists a $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ representation $\pi:\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})\rightarrow\operatorname{End}(W)$ that has the restriction
$\pi(e_{0})=\varphi_{0}X^{-1},\quad\pi(e_{1})=\varphi_{1}X,\quad\pi(t_{0})=\psi^{-1}Z^{-2},\pi(t_{1})=\psi Z^{2},$
for $\varphi_{i}$, $\psi\in\mathbb{C}^{\times}$. Then the relations $e_{i},f_{j}]=0$ for $i\neq j$ imply that
(2.19) $X\pi(f_{i})=\pi(f_{i})X,\quad i\in\{0,1\}.$
It is now useful to consider the two $U_{q}(\mathfrak{sl}_{2})$ Casimir elements $C_{i}:=qt_{i}+q^{-1}t_{i}^{-1}+(q-q^{-1})^{2}f_{i}e_{i}$. We have
$\pi(C_{0})$ $=$ $q\psi^{-1}Z^{-2}+q^{-1}\psi Z^{2}+(q-q^{-1})^{2}\varphi_{0}\,\pi(f_{0})X^{-1},$
$\pi(C_{1})$ $=$ $q\psi Z^{2}+q^{-1}\psi^{-1}Z^{-2}+(q-q^{-1})^{2}\varphi_{1}\,\pi(f_{1})X,$
and hence
(2.20) $\pi(f_{0})$ $=$ $\frac{\left(\pi(C_{0})-q\psi^{-1}Z^{-2}-q^{-1}\psi Z^{2}\right)X}{(q-q^{-1})^{2}\varphi_{0}},$
(2.21) $\pi(f_{1})$ $=$ $\frac{\left(\pi(C_{1})-q\psi Z^{2}-q^{-1}\psi^{-1}Z^{-2}\right)X^{-1}}{(q-q^{-1})^{2}\varphi_{1}}.$
The fact that $\pi(e_{i}),\pi(C_{i})]=0$ implies that $X,\pi(C_{i})]=0$. Using the fact that $ZX=qXZ$ , we see that the relations (2.20), (2.21) are incompatible with (2.19).
@@ 3.1. Factorization
For $x=z_{i}$ the equality is clear, and for $x=t_{i}$ the relation follows from the property $\chi,Z\otimes Z]=0$ and the expression for $c_{0}$. For $x=e_{0}$, the relation becomes
@@
[…]
@@
Lemma 3.6.
The polynomial ${\mathcal{P}}(Z)$ is invertible.
Proof.
This follow a very similar argument to the proof of Lemma 3.2 given in Appendix A.1. ∎
@@
$(i)\quad\mathfrak{T}_{rs}$ $:$ $\rho_{r}\otimes\bar{\rho}_{s}\rightarrow\rho_{s}\otimes\bar{\rho}_{r},$
$(ii)\quad{\mathfrak{S}}_{rs}$ $:$ $\bar{\rho}_{r}\otimes\rho_{s}\rightarrow\bar{\rho}_{s}\otimes\rho_{r}.$
@@ 3.2. Fusion
(3.3) $\pi_{z}(e_{1})=\begin{pmatrix}0&z\\ 0&0\end{pmatrix}$$\quad$ $\pi_{z}(f_{1})=\begin{pmatrix}0&0\\ z^{-1}&0\end{pmatrix}$, $\pi_{z}(t_{1})=\begin{pmatrix}q&0\\ 0&q^{-1}\end{pmatrix},\quad\pi_{z}(z_{1})=1,$
$\pi_{z}(e_{0})=\begin{pmatrix}0&0\\ z&0\end{pmatrix}$$\quad$ $\pi_{z}(f_{0})=\begin{pmatrix}0&z^{-1}\\ 0&0\end{pmatrix}$, $\pi_{z}(t_{0})=\begin{pmatrix}q^{-1}&0\\ 0&q\end{pmatrix},\quad\pi_{z}(z_{0})=1.$
@@
Let $r,s\in{\mathcal{C}}_{k}$, with $s=(x_{s},y_{s},\mu_{s})$. Defining $sq^{\pm 1}=(x_{s}q^{\pm 1},y_{s}q^{\pm 1},\mu_{s}q^{\pm 1})$, we see that we also have $sq^{\pm 1}\in{\mathcal{C}}_{k}$. Choose $z_{s}$ to satisfy $z_{s}^{2}=\kappa_{0}\kappa_{1}x_{s}y_{s}$ . Let us define the following complex constants
@@
[…]
@@
[…]
@@ 3.3. Factorization and Fusion of L-operators
where $\mathbb{I}$ is the $\times2$ unit matrix. As ${L}_{\varphi_{c}}(z)$ is idependent of both $z$ and the constant $c$, we shall henceforth denote it as simply $L_{\varphi}$.
@@
[…]
@@
[…]
@@ 3.3.2. Fusion of L-operators
[…]
@@
where $\check{\bf L}_{\Omega_{ss^{\prime}}}(z)$ is the $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ intertwiner $\pi_{z}\otimes\Omega_{ss^{\prime}}\rightarrow\Omega_{ss^{\prime}}\otimes\pi_{z}$ defined above, and the coefficients are given by
@@
[…]
@@ 4. Transfer Matrices and Q-operators
4.1. Quantum Space $V^{\otimes M}$ : the 6–Vertex Model
@@
[…]
Let us now discuss the reason for including the twist parameter $\alpha$ in this and other definitions. First, we recall that in the case of generic $q$, we are forced to introduced such a twist in order to regularise the infinite-dimensional trace involved in the definition of the Q-operator (see for example, [Kor04b]). The cyclic representations in this paper are of course $N$-dimensional, and at first sight it would appear that the twist is unnecessary. However, with cyclic representations we encounter a different issue: if we make a particular choice of the representation of the matrices $X,Z\in GL(N,\mathbb{C})$ to be $X_{ij}=\delta_{i,\hbox{mod}(j+1,N)}$, $Z_{ij}=\delta_{i,j}q^{i}$, for $i,j\in\{0,1,\cdots,N-1\}$, then it follows from basic properties of primitive roots of unity that we have
@@
where $\delta_{n\equiv 0}$ is defined to be equal to 1 if $n=0$ mod(N)(N), and 0 otherwise. Requiring $n\equiv 0$ is equivalent to ‘charge conservation mod $N$’ which is to be expected for transfer matrices associated with cyclic representations. Requiring that $m=0$ however means that the corresponding trace vanishes on all spin sectors apart from 0 mod $N$. The twisted trace is introduced in order to stop this phenomenon. Using the definition $\rho_{r}(t_{1})=q\mu_{r}Z^{2}$, we see that, provided we choose $2\alpha\notin\mathbb{Z}$, we have
@@
Thus we regularize our traces by choosing the specified twist insertion in all cases, with the requirement that $2\alpha\in\mathbb{R}\backslash\mathbb{Z}$.
@@
[…]
@@
$Q(z,\mu_{s}):$ $=$ $\operatorname*{Tr}_{W}((\rho_{s}(t_{1}^{\alpha})\otimes\mathbb{I}){L}^{M}(z,\mu_{s})\cdots{L}^{2}(z,\mu_{s}){L}^{1}(z,\mu_{s})),$
$\widebar{Q}(z,\mu_{s}):$ $=$ $\operatorname*{Tr}_{W}((\bar{\rho}_{s}(t_{1}^{\alpha})\otimes\mathbb{I})\widebar{L}^{M}(z,\mu_{s})\cdots\widebar{L}^{2}(z,\mu_{s})\widebar{L}^{1}(z,\mu_{s})),$
@@
Hence it then follows from Proposition 4.3 and the property $T(z),S_{z}]=0$ that $Q(z,\mu)$ and $\widebar{Q}(z,\mu)$ satisfy the standard TQ relations [^5] The dependence of the $Q$ operator on the parameter $\mu$, the algebraic necessity of this parameter, and an explanation how the dependence may be removed is explained in [VW20]
@@
(4.6) $T_{\Omega_{r,s}}(w)T(z_{s}/w)=C_{1}^{M}(s,w)T_{\Omega_{r,sq}}(w)+C_{2}^{M}(s,w)T_{\Omega_{r,sq^{-1}}}(w).$
@@
(which can be proved either by appealing to uniqueness of the corresponding $U_{q}(\mathfrak{b}_{+})$ intertwiner or just by checking the relation), or directly from Propositions 4.2 and 4.3 $(ii)$ plus the observation that $T_{\varphi_{c}}$ commutes with $Q_{\rho_{r}}$ and $Q_{\bar{\rho}_{s}}$. Thus we have that $T_{\Omega_{rs}}(w)$ satisfies the same relations (4.6) as the Q operator of the 6-vertex model. This fact was observed in [BS90]. One advantage of considering $T_{\Omega_{rs}}(w)$ is that one can establish its invariance under an infinite-dimensional group action known as the quantum coadjoint action. This was the approach taken in [Kor03a].
4.2. Quantum Space $W^{\otimes M}$ : the $\tau_{2}$ and Chiral Potts Models
@@
[…]
@@
$^{\prime}rrrr^{\prime}$$^{\prime}ssss^{\prime}$$^{\prime}ssss^{\prime}$$^{\prime}ssss^{\prime}$$^{\prime}ssss^{\prime}$$^{\prime}ssss^{\prime}$$^{\prime}ssss^{\prime}$${\mathcal{T}}^{CP}\sim$